# Information That Matters: Exploring Information Needs of People Affected by Algorithmic Decisions

### v1

@@ Abstract.
Explanations of AI systems rarely address the information needs of people affected by algorithmic decision-making (ADM). This gap between conveyed information and information that matters to affected stakeholders can impede understanding and adherence to regulatory frameworks such as the AI Act (Commission, 2021). To address this gap, we present the ”XAI Novice Question Bank”: A catalog of affected stakeholders’ information needs in two ADM use cases (employment prediction and health monitoring), covering the categories data, system context, system usage, and system specifications. Information needs were gathered in an interview study where participants received explanations in response to their inquiries. Participants further reported their understanding and decision confidence, showing that while confidence tended to increase after receiving explanations, participants also met understanding challenges, such as being unable to tell why their understanding felt incomplete. Explanations further influenced participants’ perceptions of the systems’ risks and benefits, which they confirmed or changed depending on the use case. When risks were perceived as high, participants expressed particular interest in explanations about intention, such as why and to what end a system was put in place. With this work, we aim to support the inclusion of affected stakeholders into explainability by contributing an overview of information and challenges relevant to them when deciding on the adoption of ADM systems. We close by summarizing our findings in a list of six key implications that inform the design of future explanations for affected stakeholder audiences.
@@ 1. Motivation
Society is in the process of negotiating how AI systems should be used in public institutions and organizations (Züger and Asghari, 2023; Lucaj et al., 2023; Lima et al., 2023). These systems should be transparent, accountable, and support human oversight, according to the legal frameworks which are being developed by regulatory bodies (Commission, 2021). But realizing these general principles in practice can be difficult as they do not always fit specific use cases. Research in the field of explainable AI (XAI) addresses this issue by examining how AI systems can be made understandable to stakeholders and how the given information enables certain actions, such as contesting a decision (Alfrink et al., 2022) or assigning accountability (Langer et al., 2021). As explanations thereby assume a key role in public AI (Züger and Asghari, 2023), the question of what constitutes a good explanation becomes central.
@@
Good explanations of AI systems are aligned with the information needs of the stakeholder audience (Miller, 2019; Mueller et al., 2019; Byrne, 2023; Ehsan et al., 2022) and are a means to help them achieve their aims, which can depend on the stakeholder’s role (Langer et al., 2021; Barredo Arrieta et al., 2020), AI literacy (Long and Magerko, 2020), and domain knowledge (Wang and Yin, 2021). However, information needs of ”affected stakeholders” – people impacted by algorithmic decision-making (ADM) (European Parliament. Directorate General for Parliamentary Research Services., 2019; Golpayegani et al., 2023) – have not yet been explored in-depth in XAI research. This can limit affected stakeholders’ options to understand ADM and, in consequence, infringe on principles of responsible and trustworthy AI (Floridi, 2019; on Artificial Intelligence, 2019; Thiebes et al., 2021; Baeza-Yate and Matthews, 2022). Concretely, providing explanations for affected stakeholders can be essential to reduce the risks of power asymmetry (Ananny and Crawford, 2018), inequality (Lopez, 2019), and informational unfairness (Schoeffer et al., 2022) in deployments of public AI (Kuziemski and Misuraca, 2020; Henman, 2020; Brown et al., 2019; de Fine Licht and de Fine Licht, 2020; Veale et al., 2018).
@@
However, building explanations to increase affected stakeholders’ understanding can be challenging: First, the group of affected stakeholders is diverse and can include ”laypeople” (Shen et al., 2020; Szymanski et al., 2021; Schmude et al., 2023a) or ”AI novices”111We use the term in reference to ”visualization novices” (Burns et al., 2023)., i.e., people with little technical literacy who have heterogeneous information needs, aims, and prior knowledge, necessitating accessible and adaptive explanations (Shulner-Tal et al., 2022; Conati et al., 2021; Kim et al., 2024). Second, the effects of explanations are influenced by perceptions of the deploying institution (Brown et al., 2019; Ehsan et al., 2021), meaning that if the institution is not trusted, the ADM system and its explanations are not trusted either (Woodruff et al., 2018; Scott et al., 2022). Third, understanding is an interconnection of knowledge and reasoning (Grimm, 2019; Keil, 2006) involving numerous cognitive processes (Keil, 2006; Lombrozo and Wilkenfeld, 2019) that can barely be captured by standardized large-scale tests (Sato et al., 2019). And lastly, AI lifecycles encompass large amounts of information (Dhanorkar et al., 2021), necessitating a preselection of what is deemed relevant.
@@
These interactions between cognitive, perceptional, and organizational dimensions can be difficult to address in explanation design. Recent XAI research, therefore, advocates for providing tailored explanations, based on the stakeholder’s aims (Dhanorkar et al., 2021; Freiesleben and König, 2023) and information needs (Liao et al., 2020, 2021), instead of explaining what is assumed to be relevant (Miller, 2023). In line with literature arguing that more research is needed on human relations with AI (Capel and Brereton, 2023), this study examines how to address the information needs of affected stakeholders and meet the challenges outlined. In this paper, we aim to support the inclusion of affected stakeholders into explainability by presenting an empirical study of their information needs, understanding, and perceptions. To this end, we use two ADM use cases as examples: an employment prediction algorithm and a health wristband for elderly care. To allow for the collection of participants’ information needs, the explanation setting is flipped: Participants independently inquire information about the systems to the end of evaluating their adoption, and are provided verbal question-driven explanations in response. We aim to address the challenges outlined above by letting participants choose independently which information they inquire, e.g., about the deploying institution or the algorithm used, reducing the amount of pre-selected information in favor of demand-based information. We analyze how the given information influences participants’ self-reported understanding and decision confidence, and how ”explanatory stances” and challenges in understanding impact participants’ information acquisition. Lastly, we analyze participants’ perceptions of risks and benefits before and after receiving explanations and describe perceived changes. Our work is thus guided by the following research questions:
RQ1-Information Needs: What are affected stakeholders’ information needs when deciding over the adoption of an ADM system?
RQ2-Understanding: How do question-driven explanations interact with participant understanding?
@@
The contributions of our work include i) a catalog of affected stakeholders’ questions titled the ”XAI Novice Question Bank” (Figure 2), in reference to Liao et al. (2020)’s XAI Question Bank; ii) recommendations on how to address understanding challenges such as blind spots and outsourcing and on how to use explanatory stances in question-driven explanations; and iii) reflections on the relevance of intentional information in explanations for affected stakeholders and considerations of how to incorporate insights from ”civic education” (Lupia, 2016) into explainability.
2. Related Work
2.1. Human-Centered Explainable AI and question-driven explanations
This work is situated in the field of explainable artificial intelligence (XAI), which aims to make algorithmic models understandable to humans (Ali et al., 2023; Barredo Arrieta et al., 2020; Adadi and Berrada, 2018; Dhanorkar et al., 2021), and, more specifically human-centred XAI (HCXAI), which examines the influence of human factors on user interaction with system explanations (Ehsan et al., 2023). Our study covers two high-risk ”algorithmic decision-making” systems (ADM systems) (European Parliament. Directorate General for Parliamentary Research Services., 2019; Commission, 2021) with humans as decision subjects, i.e., ”affected stakeholders” (Langer et al., 2021).
The term ”stakeholder” describes roles of people involved in the development or deployment of ADM, such as deployers, developers, and affected parties (Langer et al., 2021; Barredo Arrieta et al., 2020). Previous work has established the value of analyzing stakeholders’ information needs (Liao et al., 2020) to use them as basis for the design of explanations (Liao et al., 2021). As explanations of AI systems were originally used by developers to better understand the system’s behavior (Langer et al., 2021) they tend to be targeted towards technical experts, making them less suitable for ”lay people” (Shen et al., 2020; Szymanski et al., 2021; Shulner-Tal et al., 2022) or ”AI novices”. While this work thus follows the concept of question-driven explanations (Liao et al., 2021) and uses a similar approach to Liao et al. (2020)’s collection of AI practitioner’s information needs in the XAI Question Bank, the focus is shifted to another audience. In contrast to more recent work, which investigates how to adapt the XAI Question Bank to end-users (sipos2023), this work aims to provide insight on how explanations can be rendered more accessible and suitable for affected stakeholders, who are not necessarily end-users.
2.2. Perspectives on explanations and understanding
Explanations are a means to increase the recipient’s understanding of a system or decision (Langer et al., 2021), so as to enable actions such as value assessments (Kim et al., 2023), decision-making (Lee et al., 2019), and contestation (Alfrink et al., 2022). As numerous ways exist to define understanding (Grimm, 2019; Baumberger et al., 2017; Zagzebski, 2019; Schmude et al., 2023b), we use a definition that aims to preserve the epistemic aspects while allowing for operationalization in an empirical setting. We thus describe understanding as i) connecting and applying information whereas ”knowledge” only stores it (Grimm, 2019; Baumberger et al., 2017), ii) involving both sympathy and causality depending on the object that is understood (Grimm, 2019; Wiggins and McTighe, 2005), and iii) attempting to grasp the underlying structure of a phenomenon (Zagzebski, 2019).
In order to examine how explanations interact with the processes of understanding, we refer to theories by Keil (2006) and Lombrozo and Wilkenfeld (2019), stating that explanations call attention to the incompleteness of one’s understanding, prompting a reaction: (i) filling the gap with further information, (ii) ”outsourcing” it to another mind, or (iii) mentally rendering the gap irrelevant. Developing understanding, in this sense, means filling understanding gaps until one decides to have reached a ”working understanding” (Keil, 2003, 2006) and stops. The ”illusion of explanatory depth” (Rozenblit and Keil, 2002), an overestimation of one’s understanding, can perturb this process and lead to incomplete understanding (Keil, 2006; Lombrozo and Wilkenfeld, 2019; Chromik et al., 2021).
For our analysis, we further use the concept of three ”explanatory stances” (Keil, 2021, 2006; Lombrozo and Gwynne, 2014; Lombrozo and Wilkenfeld, 2019; Dennett, 1998): a theory aiming to capture how people make sense of information and generalize from it. Previous XAI literature used explanatory stances to describe how explanations are framed and processed (Páez, 2019; Zerilli, 2022; Byrne, 2023; Miller, 2019), which we use as basis for our application in an empirical setting. Following, we describe the three explanatory stances in the order mechanical, design, and intentional:
The mechanical stance222Both the terms ’mechanical’ and ’mechanistic’ are used in the literature (Keil, 2006; Lombrozo and Wilkenfeld, 2019; Keil, 2021). Where ’mechanical’ refers to the stance, ’mechanistic’ refers to the explanation. appeals to ”parts, processes, and proximate causal mechanisms” (Lombrozo and Wilkenfeld, 2019) and focuses on low-level technical detail. Explaining why an alarm clock buzzes, for example, in a mechanical stance would state that the power circuit was completed (Lombrozo and Wilkenfeld, 2019). In contrast to functional and intentional explanations, mechanistic explanations do not take into account aims or beliefs (Páez, 2019).
The design stance focuses on functions and goals beyond mechanical interactions in a system (Keil, 2006) and describes why things happen by looking at their aims and purposes (Lombrozo and Gwynne, 2014). Explaining why an alarm clock buzzes in a design stance would state that it functions according to its design: waking its owner at the set time (Lombrozo and Wilkenfeld, 2019; Dennett, 2006).
2.3. Explainability in responsible AI and regulations
Explainability is seen as a pillar of responsible AI (RAI) (Barredo Arrieta et al., 2020; Baeza-Yate and Matthews, 2022; Ehsan et al., 2023), which advocates for fair, accountable, and explainable implementations of AI systems. However, ensuring adherence to RAI principles meets practical challenges. For example, deployments of AI can lead to what Ehsan et al. (2023) and Ackerman (2000) call the sociotechnical gap: ”a divide between the technical affordances and the social needs”. This gap can lead to mismatches between what is expected of a system and what it actually does, and is further aggravated if stakeholders’ interests or values are in conflict (Jakesch et al., 2022).
The draft AI Act, an EU regulation aiming to guide the development and deployment of high-risk AI systems, prescribes measures to uphold RAI principles (Commission, 2021; Panigutti et al., 2023; Sovrano et al., 2022; Veale and Zuiderveen Borgesius, 2021), such as ”the provision of relevant documentation and information to the user” (Panigutti et al., 2023). Which information is considered relevant and how it should be conveyed, however, remains unspecified (Nannini et al., 2023). Lucaj et al. (2023) therefore argue that doubts remain as to how the high-level regulations can improve explainability in practice. To ensure that information is both relevant and intelligible for affected stakeholders, their information needs ought to be elicited and incorporated (Kim et al., 2023). This work aims to contribute to this process by giving an overview of affected stakeholders’ information needs, to be used for explanation design (Madaio et al., 2020; Hagendorff, 2022; Ferrario and Loi, 2022).

### v6

@@ Abstract.
Every AI system that makes decisions about people has a group of stakeholders that are personally affected by these decisions. However, explanations of AI systems rarely address the information needs of this stakeholder group, who often are AI novices. This creates a gap between conveyed information and information that matters to those who are impacted by the system’s decisions, such as domain experts and decision subjects. To address this, we present the “XAI Novice Question Bank,” an extension of the XAI Question Bank (Liao et al., 2020) containing a catalog of information needs from AI novices in two use cases: employment prediction and health monitoring. The catalog covers the categories of data, system context, system usage, and system specifications. We gathered information needs through task-based interviews where participants asked questions about two AI systems to decide on their adoption and received verbal explanations in response. Our analysis showed that participants’ confidence increased after receiving explanations but that their understanding faced challenges. These included difficulties in locating information and in assessing their own understanding, as well as attempts to outsource understanding. Additionally, participants’ prior perceptions of the systems’ risks and benefits influenced their information needs. Participants who perceived high risks sought explanations about the intentions behind a system’s deployment, while those who perceived low risks rather asked about the system’s operation. Our work aims to support the inclusion of AI novices in explainability efforts by highlighting their information needs, aims, and challenges. We summarize our findings as five key implications that can inform the design of future explanations for lay stakeholder audiences.
@@ 1. Motivation
Society is in the process of negotiating which AI systems can be used responsibly in public institutions and organizations (Züger and Asghari, 2023; Lucaj et al., 2023; Lima et al., 2023). These systems should be transparent, accountable, and support human oversight, according to the legal frameworks that are being developed by regulatory bodies (European Commission, 2021). Research in the field of explainable AI (XAI) addresses the realization of these values by examining how AI systems can be made understandable to stakeholders and how the given information enables specific actions, such as the contestation of decisions (Alfrink et al., 2023) and the assignment of accountability (Langer et al., 2021). As explanations thereby assume a key role in the deployment of responsible AI systems for the public (Züger and Asghari, 2023), the question of what constitutes a good explanation becomes central.
@@
Good explanations of AI systems are aligned with the information needs of the stakeholders (Miller, 2019; Mueller et al., 2019; Byrne, 2023; Ehsan et al., 2022) and are a means to help them achieve their aims, which can depend on the stakeholder’s role (Langer et al., 2021; Barredo Arrieta et al., 2020), AI literacy (Long and Magerko, 2020), and domain knowledge (Wang and Yin, 2021). However, information needs of people who are impacted directly and personally by algorithmic decision-making (ADM) systems (European Parliament. Directorate General for Parliamentary Research Services., 2019; Golpayegani et al., 2023), such as domain expert users and decision subjects, have not yet been explored in-depth in XAI research. As a result, current explanations rarely meet the needs of these stakeholder groups. This is critical, as providing explanations for these stakeholder groups can be essential to reduce the risks of power asymmetry (Ananny and Crawford, 2018), inequality (Lopez, 2019), and informational unfairness (Schoeffer et al., 2022) in deployments of public AI (Kuziemski and Misuraca, 2020; Henman, 2020; Brown et al., 2019).
@@
However, building explanations that aid domain experts and decision subjects in understanding ADM systems can be challenging. First, these stakeholder groups are diverse and can include “AI novices”111We use the term in reference to “visualization novices” (Burns et al., 2023) and synonymously with “lay people” (Shen et al., 2020; Szymanski et al., 2021; Schmude et al., 2023)., i.e., people with little technical literacy who have heterogeneous information needs and prior knowledge, which necessitates accessible and adaptive explanations (Shulner-Tal et al., 2022; Conati et al., 2021; Kim et al., 2024). Second, the effects of explanations are influenced by perceptions of the deploying institution (Brown et al., 2019; Ehsan et al., 2021), meaning that if the institution is not trusted, the ADM system and its explanations are not trusted either (Woodruff et al., 2018; Scott et al., 2022). Third, understanding is an interconnection of knowledge and reasoning (Grimm, 2019; Keil, 2006) involving numerous cognitive processes (Keil, 2006; Lombrozo and Wilkenfeld, 2019) that can barely be captured by standardized large-scale tests (Sato et al., 2019). Lastly, AI lifecycles encompass large amounts of information (Dhanorkar et al., 2021) and, therefore, require a preselection of what is deemed relevant.
These interactions between cognitive, perceptional, and organizational dimensions can be difficult to address in explanation design. Recent XAI research, therefore, advocates for providing tailored explanations, based on the stakeholder’s aims (Dhanorkar et al., 2021; Freiesleben and König, 2023) and information needs (Liao et al., 2020, 2021), instead of explaining what is assumed to be relevant (Miller, 2023). This study thus follows research that aims to build “human-centered” explanations of AI by making them more accessible and better suited to human cognitive processes (Shin, 2023; Ehsan et al., 2022).
@@
This paper contributes to including domain experts and decision subjects in explainability by presenting an empirical study of their information needs, understanding, and perceptions. We use two ADM use cases as examples: an employment prediction algorithm and a health wristband for geriatric care. To collect participants’ information needs, the explanation setting is flipped: Participants ask questions about the algorithmic systems to evaluate their adoption and are given verbal explanations in response. Participants thus independently choose which information they inquire about, reducing the amount of pre-selected information in favor of demand-based information. We analyze the impact of question-driven explanations on participants’ self-reported understanding and decision confidence, as well as the role of “explanatory stances” (Keil, 2006) in their information acquisition. Lastly, we analyze participants’ perceptions of risks and benefits before and after receiving explanations and describe perceived changes. Our work is guided by the following research questions:
[RQ1]
Information Needs: What information do AI novices who could be affected by algorithmic decisions need in order to decide about adopting an ADM system?
[RQ2]
Understanding: How do question-driven explanations interact with participants’ understanding?
[RQ3]
Perception: How do question-driven explanations impact perceptions of the systems’ risks and benefits?
The contributions of our work include i) a catalog of affected stakeholders’ information needs summarized in the “XAI Novice Question Bank” (Figure 3)222In reference to Liao et al. (2020)’s XAI Question Bank.; ii) recommendations on how to address understanding challenges such as blind spots and outsourcing and on how to use explanatory stances in question-driven explanations; and iii) reflections on the relevance of information about intention in explanations for affected stakeholders and considerations of how to incorporate insights from “civic education” (Lupia, 2016) into explainability.
2. Background and related work
We structured our literature review in three sections. The first section describes related work in Human-Centered AI on high-risk systems and stakeholder-focused explainability (Section 2.1). The second section covers the interplay between explanations and understanding as well as definitions of understanding and approaches to analyze it (Section 2.2). The third section describes the importance of people’s perceptions of ADM systems’ risks and benefits and the relationship between explanations and the assessment of normative values (Section 2.3).
2.1. Human-Centered AI (HCAI)
The notion of human-centered AI aims to provide principles and frameworks that enable “ethical, interactive, and contestable use” of AI systems, covering both interaction design approaches as well as the incorporation of ethical values (Capel and Brereton, 2023). Shneiderman (2022), in line with other scholars (Araujo et al., 2020; Xu, 2019; Shin, 2023; Shin and Shin, 2023), emphasizes that human-centered AI systems should 1) build on user observation and stakeholder engagement and should 2) empower rather than replace people, i.e., be “accessible and controllable” Shin and Shin (2023). Prioritizing people, in this sense, means giving attention to users and other stakeholders throughout the development, deployment, and evaluation of AI systems (Shin, 2023) while observing human values such as fairness, trust, and accountability (Shneiderman, 2022). Similar notions have been brought forward in multiple responsible and trustworthy AI guidelines, such as the Montreal Declaration for Responsible Development of AI, the Vienna Manifesto on Digital Humanism, and, more recently, the EU AI Act (European Commission, 2021).
Human-centered design becomes especially important when considering “high-risk” AI applications – systems that pose risks to the health, safety, or fundamental rights of persons (European Commission, 2021). Rendering high-risk systems transparent, understandable, and explainable to users and preserving their control is seen as a mitigation of these risks (Adadi and Berrada, 2018), but the practical realization proves challenging (Ananny and Crawford, 2018). As people who are affected by ADM systems tend to be AI novices, explanation approaches need to be tailored to their needs, but as Shin (2023) notes, there are few “well-established ways of incorporating lay users into the process of designing algorithms”. Further, no universal explanation method exists that can cover the needs of all stakeholder groups equally (Phillips et al., 2021). Therefore, the question of how AI novices’ information needs can be analyzed and incorporated into explanation approaches is an open challenge in XAI research (Shulner-Tal et al., 2022).
To this end, several studies have examined how explanations can help lay users to understand algorithmic systems: Szymanski et al. (2021) examined explanations for model predictions and found that lay users preferred a visual format over a textual one but had difficulties correctly interpreting the visual explanation, which led to the development of a new explanation combining both formats. Cheng et al. (2019) found that interactive explanations and white-box models improved lay participants’ comprehension; however, a later study by Bove et al. (2022) could not replicate the results regarding interactive explanations and found the adverse effect. Bove et al. (2022) suggest that the diverging results could have stemmed from the difference in task domains (student admission vs. car insurance) and emphasize that context influences the explanation’s effect. Lastly, Bertrand et al. (2023) used interviews to elicit the information needs of domain experts and end-users about a financial robo-advisor, which they then used to develop and test explanations for the system’s recommendations in a large-scale survey. Their analysis showed that feature-based SHAP explanations, one of the most well-known explanation methods, did not increase users’ understanding nor enable a suitable trust calibration. In contrast, our study takes a more general perspective on the qualitative exploration of participants’ global information needs and perceptions when they vote on the adoption of a high-risk ADM system.
These multi-faceted insights indicate how difficult it can be to tailor explanations of algorithmic systems to lay people. For this reason, this study aims to take a closer look at the information needs of AI novices and how they can be categorized. To elicit these needs, participants in our study assume the role of a decision-maker who can ask questions and receive explanations in return, as described in Section 3. We thus “flip” the explanation setting to answer the following research question:
RQ1-Information Needs: What information do AI novices who could be affected by algorithmic decisions need in order to decide about adopting an ADM system?
2.2. Explanations and understanding
One purpose of explanations is to increase the recipient’s understanding of a system (Langer et al., 2021). This increased understanding can help people to pursue their own goals, such as to assess the system’s values (Kim et al., 2023), consider its recommendations (Lee et al., 2019), or contest its decisions (Alfrink et al., 2023). Shin et al. (2022a) further state that the understandability of explanations impacts how users evaluate both the quality of explanations and the system’s fairness, accountability, and transparency. However, despite this vital role of understanding, there is no general conception of what is most important about an AI system, nor how this understanding can be evaluated (Schmude et al., 2023).
Numerous definitions of understanding have been proposed in the literature (Grimm, 2019; Baumberger et al., 2017; Zagzebski, 2019; Keil, 2006). For this work, we use a definition that aims to preserve its epistemic aspects while allowing for operationalization in an empirical setting. We thus define understanding as i) connecting and applying information, whereas “knowledge” only stores information (Grimm, 2019; Baumberger et al., 2017), ii) involving different varieties depending on whether people (intention, norms) or objects (causality, natural laws) are understood (Grimm, 2019), and iii) being the attempt to grasp the underlying structure of a phenomenon by way of simplification (Zagzebski, 2019).
2.2.1. Gaps in understanding and explanatory stances
To integrate the epistemic aspects of understanding practically, we use models provided by the cognitive sciences. Keil (2006) and Lombrozo and Wilkenfeld (2019) state that explanations increase understanding by calling attention to gaps in understanding. These gaps can be dealt with by (i) filling them with new information, (ii) “outsourcing” them to another mind, or (iii) deciding that they are irrelevant. In this sense, developing understanding means dealing with understanding gaps until one feels to have reached a “working understanding” (Keil, 2003, 2006). This process can be perturbed by the “illusion of explanatory depth” (Rozenblit and Keil, 2002), an overestimation of one’s understanding, which can lead to incomplete understanding (Keil, 2006; Lombrozo and Wilkenfeld, 2019; Chromik et al., 2021).
For our analysis, we further use the concept of three explanatory stances (Keil, 2021, 2006; Lombrozo and Gwynne, 2014; Lombrozo and Wilkenfeld, 2019), introduced by Dennett (1998): a theory aiming to capture how people make sense of information and generalize from it in order to develop predictive strategies. The three explanatory stances are described below in the order mechanical, design, and intentional stance.
The mechanical stance focuses on low-level technical detail, such as “parts, processes, and proximate causal mechanisms” (Lombrozo and Wilkenfeld, 2019). For example, explaining why an alarm clock buzzes in a mechanical stance would state that the circuit connecting the power source and buzzer was completed (Lombrozo and Wilkenfeld, 2019). In contrast to functional and intentional explanations, mechanistic333Both the terms ’mechanical’ and ’mechanistic’ are used in the literature (Keil, 2006; Lombrozo and Wilkenfeld, 2019; Keil, 2021). Where ’mechanical’ refers to the stance, ’mechanistic’ refers to the explanation. explanations do not take into account aims or beliefs (Páez, 2019).
The design stance focuses on functions and goals beyond mechanical interactions in a system (Keil, 2006) and describes why things happen by looking at their functional purposes (Lombrozo and Gwynne, 2014). Explaining why an alarm clock buzzes from a design stance would state that it functions according to its design: waking its owner at the set time (Lombrozo and Wilkenfeld, 2019; Dennett, 2006).
The intentional stance focuses on a system’s purpose as defined by the human beliefs or aims ingrained in it (Keil, 2006; Dennett, 1998; Yurrita et al., 2022). Explaining why an alarm clock buzzes in an intentional stance might refer to its purpose of waking its owner, whose desire is to have enough time before leaving for work.
Previous XAI literature used explanatory stances to examine how explanations are perceived and evaluated (Páez, 2019; Zerilli, 2022; Miller, 2019; Byrne, 2023). For example, Byrne (2023) describes that explanatory stances shed light on how people explain the behavior of algorithmic systems to themselves, which might differ depending on whether they interpret the decisions as results of an algorithmic process or as depictions of decisions by a human (e.g., a job counselor). The stances, therefore, aid in identifying whether the users’ information needs focus on the systems’ technical or intentional aspects and thus make it easier to match their information needs to a specific explanation (Byrne, 2023). In our approach, we use the stances to differentiate if participants asked for information about technical details, functional relations, or the sociotechnical context.
2.2.2. Question-driven explanations
Questions assume a central role in XAI, as explanations of AI systems can be defined as responses to questions that are asked by the user. Miller (2019) defines three classes of these questions (what, how, and why), which are meant to cover different levels of causal reasoning through their responses. A comparable classification approach has been developed earlier by Lim and Dey (2009), who defined five intelligibility questions (what, why, why not, what if, and how to) to capture questions that “end-users of novel systems may ask” to convey to users the practical functionality of these systems. A practical implementation of question-driven explanation was pursued even earlier by Ram (1989), who built AQUA – a system that used questions about a story’s content to generate explanations and advance its knowledge.
From a didactic perspective, Wiggins and McTighe (2005) describe questions to be “doorways to understanding”, through which learners explore concepts, theories, and problems that reside within the content. When we thus define explanations as answers to questions, it is noteworthy that these questions often remain tacit in XAI. In response, Liao et al. (2021) propose a question-driven design process for explanations, which is grounded in the needs and questions of users. The concept further builds upon their previous work on the XAI Question Bank (Liao et al., 2020), a collection of expert users’ information needs represented as prototypical questions (e.g., “What kind of output does the system give?”). In our setting, we apply the question-driven design process by providing participants with explanations as responses to their questions. We aim to examine how participants’ understanding is changed by these explanations and whether they allow for a better matching between information and explanatory stances. To this end, we pose the following research question:
RQ2-Understanding: How do question-driven explanations interact with participants’ understanding?
2.3. Perceptions of risks and benefits of algorithmic systems
The acceptance and responsible use of algorithmic systems in society majorly depends on how these systems are perceived (Jakesch et al., 2022; Shin, 2022; Züger and Asghari, 2023). If the “technical affordances” of a system diverge from the “social needs” (Ehsan et al., 2023), i.e., if the system does not or cannot do what is expected from it, it can result in users’ distrust in the deploying institution and their rejection of the system (Jakesch et al., 2022). Explainability is understood to influence these perceptions, taking a “facilitating” role in how users evaluate an algorithm’s features and their attitudes towards it (Shin, 2021; Langer et al., 2021). Shin (2023) describes this as a “dual-step flow”, in which explainability allows users to determine their perceptions of the system’s fairness, transparency, and accountability (non-functional attributes), which then, in turn, lets them calibrate their trust in the system and their evaluation of accuracy and personalization (functional attributes). Trust, in this sense, takes a mediating role in that it connects explainability to the system’s evaluation by the user (Shin, 2022).
Explanations thus play an essential role in helping people adjust their perceptions of whether an algorithmic system adheres to normative values. But to fulfill this facilitating function, the information provided in explanations needs to match “what users really wish to understand” (Shin, 2023). Ensuring this fit can further improve the perceived “informational fairness” of algorithmic systems, meaning that information is consistent, reveals relationships between input and output, and includes actionable insights (Schoeffer et al., 2022). However, several studies have shown the complex interrelations that exist between normative values, trust, acceptance, and explanations (Shulner-Tal et al., 2022; Shin et al., 2022b; Schoeffer et al., 2022), and the meaning of normative values such as fairness, trust, and transparency is not strictly fixated in colloquial speech, making it difficult to compare self-reports by participants (Schmude et al., 2023). Motivated by Corvite et al. (2023)’s work, we therefore use an evaluation based on perceived risks and benefits of algorithmic systems. By using a two-dimensional scale that depicts both risk and benefit simultaneously (described in Section 3), we further aim to let participants consider the positive and negative consequences of deploying a specific algorithmic system. As we focus on the explanations’ effect on these perceptions, we formulate the third research question: